OpenStudy (anonymous):

Simlify the expression: 80^(1/4)/5^(-1/4) Simlify the expression: 80^(1/4)/5^(-1/4) @Mathematics

OpenStudy (agreene):

\[\frac{80^{1/4}}{5^{-1/4}}\] is that the question?

OpenStudy (anonymous):

Yes!!!

OpenStudy (agreene):

\[\frac{80^{1/4}}{5^{-1/4}}=\frac{2*5^{1/4}}{\frac{1}{5^{1/4}}}=2\sqrt 5\]

OpenStudy (anonymous):

Thanks!